It takes 150 J of energy to heat 10 g of substance 1 from 25oC to 50oC. It takes 250 J of energy to heat 10 g of substance 2 from 25oC to 50oC. Which substance has a higher specific heat capacity?
1
Expert's answer
2020-05-08T14:11:52-0400

Substance 1: c=Q/(m*ΔT)=150/(10*25)=0.6 J/goC

Substance 2: c=Q/(m*ΔT)=250/(10*25)=1 J/goC

Answer: Substance 2 has a higher specific heat capacity